# Numaan NaeemKaushal Kumar MauryaKseniia PetukhovaEkaterina Kochmar

### v1

@@ Appendix A Pedagogical Dimensions and Data
A.1 The Task
The proposed evaluation toolkit targets the student mistake remediation (SMR) task within educational dialogues between a student and a tutor in the mathematical domain. Formally, let the conversation history be defined as $H=\{(T_{1},S_{1}),(T_{2},S_{2}),\ldots,(T_{t},S_{t})\},$ where $T_{i}$ is the tutor’s $i$-th response and $S_{i}$ is the student’s $i$-th response. Let $S_{k}$ denote the student’s most recent $k$ utterances, where $k\in[1,\ldots,t]$, that contain a mistake or confusion. The tutor’s objective is to generate the next response $T_{t+1}$ that best addresses this mistake or confusion. The backend module of the toolkit focuses on evaluating the pedagogical soundness of the $T_{t+1}$ response.
A.2 Evaluation Dimensions
@@
Mistake identification Has the tutor identified a mistake in a student’s response?
(1) Yes
(2) To some extent
(3) No
Yes
Mistake location Does the tutor’s response accurately point to a genuine mistake and its location?
(1) Yes
(2) To some extent
(3) No
Yes
Providing guidance
Does the tutor offer correct and relevant guidance, such as an explanation,
elaboration, hint, examples, and so on?
(1) Yes
(2) To some extent
(3) No
Yes
Actionability Is it clear from the tutor’s feedback what the student should do next?
(1) Yes
(2) To some extent
(3) No
Yes
@@
A.3 Dataset Details and Statistics
@@ B.1 Motivation
Table 4: Comparison of macro-F1 scores and model parameters between the our automated evaluation model (i.e., LoMTL) and the top-performing teams in the BEA shared task across four dimensions (aka. tasks). Note that, for a fair comparison, we include only the teams that participated in all four tracks of the shared task.
@@ B.3 Prompts and Configurations
Appendix C Evaluation Experiment Details
C.1 Metrics Across Dimensions
The per-dimension results, presented in Table˜6 show that Prometheus2 performs poorly, while our model notably outperforms GPT-5 on the Mistake Identification and Actionability dimensions. However, it underperforms GPT-5 by 3 and 6 percentage points in terms of macro-F1 for Mistake Location and Providing Guidance, respectively. A similar trend is observed in the ten dialogues used for the demonstration.
@@ C.1 Metrics Across Dimensions
Accuracy Macro-F1 Accuracy Macro-F1
MI ML PG AC MI ML PG AC MI ML PG AC MI ML PG AC
C.2 Full Questionnaire
@@ Appendix D User Interface (UI) Details
Figure 12: Visualization of the Tutor Performance Summary panel in the Visualizer module, displaying the aggregated evaluation scores for each tutor across all assessment dimensions within the MRBench development dataset.
Figure 13: Interface of the Visualization Controls Panel within the Visualizer module, showing configurable options for selecting tutors and evaluation dimensions to generate comparative spider chart and bar plot visualizations.
Figure 14: Visualization of the Dataset Visualization panel within the Visualizer module, showcasing the spider chart representation of tutor performance across selected evaluation dimensions, based on configurations chosen in the Visualization Controls Panel.
Figure 15: Visualization of the Dataset Visualization panel within the Visualizer module, showcasing the bar plot representation of tutor performance across selected evaluation dimension, based on configurations chosen from the Visualization Controls Panel.

### v2

@@ Appendix A Pedagogical Dimensions and Data
A.1 Evaluation Dimensions
@@
Mistake identification Has the tutor identified a mistake in a student’s response? (1) Yes (2) To some extent (3) No Yes
Mistake location Does the tutor’s response accurately point to a genuine mistake and its location? (1) Yes (2) To some extent (3) No Yes
Providing guidance Does the tutor offer correct and relevant guidance, such as an explanation, elaboration, hint, examples, and so on? (1) Yes (2) To some extent (3) No Yes
Actionability Is it clear from the tutor’s feedback what the student should do next? (1) Yes (2) To some extent (3) No Yes
@@
A.2 Dataset Details and Statistics
@@ B.1 Motivation
Prometheus2 Kim et al. (2024) 0.410 4 (prometheus-7b-v2.0) $4\times 7\text{B}=28\text{B}$ $14\times$ larger
GPT-5 Achiam et al. (2023) 0.581 - - -
@@
Table 5: Comparison of macro-F1 scores and model parameters between our automated evaluation model (i.e., LoMTL) and the top-performing teams in the BEA shared task and LLM-as-a-judge models across four dimensions (aka. tasks). Note that, (1) for a fair comparison, we include only the teams that participated in all four tracks of the shared task, and (2) since GPT-5 is a closed-source model, its parameter details are not publicly available.
@@ B.3 Prompts and Configurations
Appendix C Toolkit Evaluation Details and Results
C.1 Extended Evaluation Results
In addition to the observations on the accuracy and macro-F1 scores in Section 4.1, a closer inspection of precision and recall (from Table 7) further supports our findings. On the full test set, LoMTL achieves the highest macro-precision (0.63) while maintaining competitive recall (0.59), indicating more accurate and consistent positive predictions compared to both baselines. GPT-5 attains slightly higher recall (0.60) but with lower precision (0.58), suggesting a more recall-oriented behavior. On the demonstration subset, GPT-5 achieves the highest precision (0.60) and recall (0.61), whereas LoMTL remains competitive (0.58 precision, 0.56 recall) and substantially outperforms Prometheus2. Overall, these results show that LoMTL maintains a balanced precision–recall trade-off, reinforcing its robustness across evaluation dimensions.
C.2 Human Annotation Analysis
@@
Accuracy Macro-F1 Precision Recall Accuracy Macro-F1 Precision Recall
LoMTL (ours) 0.72 0.60 0.63 0.59 0.68 0.55 0.58 0.56
Prometheus2 0.47 0.41 0.44 0.45 0.41 0.34 0.38 0.36
GPT-5 0.66 0.58 0.58 0.60 0.66 0.59 0.60 0.61
Table 7: Accuracy, macro-F1, macro-precision and macro-recall scores (averaged across Mistake Identification (MI), Mistake Location (ML), Providing Guidance (PG), and Actionability (AC)) for our LoMTL model, Prometheus2, and GPT-5 on the full test set from Kochmar et al. (2025) and on the demonstration set. Best results are shown in bold.
C.3 Full Questionnaire
@@ Appendix D User Interface (UI) Details
Figure 12: Tutor Performance Summary panel in the Visualizer module, displaying the aggregated evaluation scores for each tutor across all assessment dimensions within the MRBench development dataset.
Figure 13: Interface of the Visualization Controls Panel, showing configurable options for selecting tutors and evaluation dimensions to generate comparative spider chart and bar plot visualizations.
Figure 14: The spider chart representing tutors performance across selected evaluation dimensions, based on configurations chosen in the Visualization Controls Panel.
Figure 15: The bar plot representing tutors performance across selected evaluation dimension, based on configurations chosen from the Visualization Controls Panel.